OpenStudy (anonymous):

Which is not a factor of 48? A. 8 B. 14 C. 16 D. 24

OpenStudy (anonymous):

which of these number is not divisible by 48

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

48/8=? 48/14= 48/16= 48/24=